OpenStudy (anonymous):

Use an inequality and the five step process to solve each problem The women’s volleyball team can spend at most $700.00 for its awards banquet at a local restaurant. If the restaurant charges a $100.00 set up-fee plus $24 per person, at most how many can attend?

OpenStudy (watchmath):

Let x be the number of person, then for these x person the team needs to spent \(1000+24x\) Since the available budget is $700, then \(100+24x\leq 700\) \(24x\leq 600\) \(x\leq 25\) So at most 25 person can attend the banquet

OpenStudy (immanuelv):

25 can only attend!

OpenStudy (anonymous):

tks guys